OpenStudy (anonymous):

Integration! \[\int x^2\sqrt{x-2} dx\]

OpenStudy (anonymous):

So I think I have to \(u = x - 2\) However that leads to \(u + 2 = x\) \[\int (u+2)^2\sqrt{u} du\] I just wonder if anyone had a faster way before I dove headlong into that mess. :D

OpenStudy (primeralph):

It works doesn't it?

OpenStudy (anonymous):

Well, yes. :D

OpenStudy (primeralph):

So don't sweat it.

OpenStudy (anonymous):

\[\int (u^2 + 4u + 4)\sqrt{u}\]\[\int u^{5/2} + 4u^{3/2} + 4u^{1/2}\]

OpenStudy (anonymous):

\[\frac{2u^{7/2}}{7} + \frac{8u^{5/2}}{5} + \frac{8u^{3/2}}{3} + C \] \[\frac{2(x-2)^{7/2}}{7} + \frac{8(x-2)^{5/2}}{5} + \frac{8(x-2)^{3/2}}{3} + C \]

OpenStudy (primeralph):

Oh well, I guess there's another way, but, there you go.

OpenStudy (anonymous):

What's the other way? :O

OpenStudy (dan815):

i remember seeing a question like this

OpenStudy (primeralph):

You asked for it.

OpenStudy (primeralph):

Forget it; it's much longer. Stick with that.

OpenStudy (anonymous):

Thanks!

OpenStudy (primeralph):

It's all you. Good job.